OpenStudy (anonymous):

Two forces with magnitudes of 200 and 100 pounds act on an object at angles of 60° and 170° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

Change the forces into Cartesian coordinates, and add them using vector addition. You can then change back if you want.

OpenStudy (anonymous):

Changing to Cartesian from Polar of magnitude \(r\) and angle \(theta\): \[ x = r\cos\theta\\ y = r\sin\theta \]Changing from Polar to Cartesian of horizontal magnitude \(x\) and vertical magnitude \(y\): \[ r = \sqrt{x^2+y^2} \\ \theta = \tan^{-1}(y/x) \]

OpenStudy (anonymous):

x=200cos60 ? and y= 100sin170?

OpenStudy (anonymous):

\[ F_1 = 200\cos60^\circ\hat{\imath} + 200\sin60^\circ\hat{\jmath}\\ F_2 = 100\cos170^\circ\hat{\imath} + 100\sin170^\circ\hat{\jmath} \]

OpenStudy (anonymous):

\[ F_3 = F_1+F_2 \]

OpenStudy (anonymous):

Is F3 my my direction or magnitude? and F3=100+100 -98.48-98.48?

OpenStudy (anonymous):

.... You really don't understand how vectors work?

OpenStudy (anonymous):

\(F_3\) is a vector.

OpenStudy (anonymous):

So are \(F_1\) and \(F_2\)... do you know how to add vectors?

OpenStudy (anonymous):

If i had <1,5> + <3,8> = <4,13> is that what you mean?

OpenStudy (anonymous):

Yeah... That's it. Find \(F_1\) and \(F_2\) first.

OpenStudy (anonymous):

I dont understand why I have to do F1=200cos60∘ı^+200sin60∘ȷ^ instead of just F1=200cos60∘ı^ and also what is ∘ı^? Is it just supposed to be degrees?

OpenStudy (anonymous):

It's vector notation...

OpenStudy (anonymous):

\(<3,2> = 3\hat{\imath}+2\hat{\jmath}\)

OpenStudy (anonymous):

Nevermind lol! That was a silly question, i didnt look at the fact that cos for the first and then sin :)

OpenStudy (anonymous):

F1=200cos60,200sin60 =100,173.2 F2=100cos170,100sin170 =-98.48,17.36

OpenStudy (anonymous):

F3=1.52,190.56

OpenStudy (anonymous):

Is this the end result?

OpenStudy (anonymous):

For the direction using θ=tan−1(y/x) which numbers would i plug in for x and y?

OpenStudy (anonymous):

\[\tan^{-1} (170/60) or \tan^{-1} (60/170)\]

OpenStudy (anonymous):

Ummm, well in this case \(x =1.52\) and \(y=190.56\)

OpenStudy (anonymous):

\[ \theta = \tan^{-1}(190.56/1.52) \]

OpenStudy (anonymous):

\[ r = \sqrt{(1.52)^2+(190.56)^2} \]